# A Bayesian likely responder approach for the analysis of randomized controlled trials

### v1

@@ B. Simulation-Based Sensitivity Analysis
B3. Sensitivity to Correlation between $E[Y|T,x]$ and E[Y|T,x]E[Y|C,x]E[Y|T,x]-E[Y|C,x]
We simulated a data generative mechanism in which correlation between $E[Y|T,x]$ and E[Y|T,x]E[Y|C,x]E[Y|T,x]-E[Y|C,x] was close to 1. Results for the proposed Bayesian two-stage model are shown in Table S3. The corrected estimator continued to yield accurate and well-calibrated subgroup effect estimates.
Table S3: Simulation results for both binary and continuous outcomes under a data generative mechanism where prognosis and treatment effect are nearly perfectly correlated.
@@
375 (125) UR 0.174 0.153 -0.145 0.499 0.960 0.124 0.114 -0.101 0.371 0.960
LR 0.173 0.164 0.095 0.493 0.965 0.111 0.104 0.087 0.368 0.990
750 (250) UR 0.091 0.084 -0.082 0.334 0.965 0.055 0.053 -0.049 0.249 0.950
LR 0.081 0.068 0.114 0.334 0.975 0.065 0.059 0.075 0.249 0.945
1500 (500) UR 0.048 0.039 -0.093 0.234 0.960 0.033 0.032 -0.035 0.170 0.920
LR 0.052 0.039 0.115 0.232 0.940 0.030 0.029 0.020 0.171 0.960
@@ D. Variable Importance in the BART Model for COMPILE COVID-19 Clinical Trial
In this section, we report results supplementary to Section 5.2 of the main manuscript, regarding the prognostic balancing score model developed in the design stage. In the context of BART, variable importance is conveyed through the Variable Inclusion Proportion (VIP).[bleich2014variable] The VIP represents the proportion of posterior MCMC samples in which a variable was included in a splitting rule in the sum-of-trees model, as described by Chipman et al.[chipman2010bart]. Figure S1 displays the VIPs from a BART model trained on all treated individuals. In order to sharpen our variable selection, we limited the model to 20 trees when computing the VIPs.
@@
Patients with preexisting conditions—cardiovascular disease, pulmonary disease, and diabetes—were, on average, predicted to have a lower probability of the outcome event compared to those without these conditions, despite all receiving CCP treatment. These findings align with results from the multivariate model reported by Wu et al.[wu2024bayesian], which suggested that CCP treatment was more effective for patients with preexisting disease, given the reference levels of other pre-treatment characteristics. However, as shown in Figure 2 in main text, the contribution of these effects was minimal, making their clinical significance uncertain.

### v2

@@ B. Simulation-Based Sensitivity Analysis
B3. Sensitivity to Alignment between $E[Y\mid T,x]$ and $E[Y\mid T,x]-E[Y\mid C,x]$
The likely responder (LR) subgroup is defined by a clinically prespecified threshold on $s(x)=E[Y\mid T=1,x]$ and, in general, does not coincide with selection based on the individual treatment effect $\Delta(x)=E[Y\mid T=1,x]-E[Y\mid T=0,x]$. To examine the behavior of the proposed two-stage Bayesian procedure under different relationships between prognosis and treatment effect, we conducted sensitivity analyses under data-generating mechanisms with varying degrees of alignment between $s(x)$ and $\Delta(x)$.
Across all scenarios, covariates were generated as $X_{i}\sim N(0,I_{10})$, and the LR subgroup was defined using a common clinical criterion with $\texttt{minCond}=0$, corresponding to $s(x)\geq 0$.
For the binary outcome, responses were generated from $\text{logit}\{E(Y\mid T,X)\}=\boldsymbol{\mu}_{0}^{\top}X+\boldsymbol{\mu}_{1}^{\top}X\cdot T$.
For the continuous outcome, responses were generated from $E(Y\mid T,X)=\boldsymbol{\mu}_{0}^{\top}X+\boldsymbol{\mu}_{1}^{\top}X\cdot T$.
Different choices of $(\boldsymbol{\mu}_{0},\boldsymbol{\mu}_{1})$ were used to induce varying degrees of alignment between $s(x)$ and $\Delta(x)$, as described below.
Strong positive alignment.
To induce strong positive alignment between $s(x)$ and $\Delta(x)$ (Table S3), we set $\boldsymbol{\mu}_{0}=\begin{bmatrix}1.2,\,1.0,\,0.8,\,0.6,\,0.4,\,0.2,\,-0.2,\,-0.4,\,-0.6,\,-0.8\end{bmatrix}^{\top}$,
with $\boldsymbol{\mu}_{1}=\begin{bmatrix}1.08,\,0.90,\,0.72,\,0.54,\,0.36,\,0.18,\,-0.18,\,-0.36,\,-0.54,\,-0.72\end{bmatrix}^{\top}\quad\text{(binary outcome)}$,
and $\boldsymbol{\mu}_{1}=\begin{bmatrix}0.36,\,0.30,\,0.24,\,0.18,\,0.12,\,0.06,\,-0.06,\,-0.12,\,-0.18,\,-0.24\end{bmatrix}^{\top}\quad\text{(continuous outcome)}$.
Weak or moderately negative alignment.
To induce weak or moderately negative alignment between $s(x)$ and $\Delta(x)$ (Table S4), we retained the same $\boldsymbol{\mu}_{0}$ and set
$\boldsymbol{\mu}_{1}=\begin{bmatrix}0.020,\,-0.255,\,-0.095,\,-0.015,\,0.081,\,-0.002,\,0.115,\,-0.109,\,0.000,\,0.203\end{bmatrix}^{\top}\quad\text{(binary outcome)}$, and
$\boldsymbol{\mu}_{1}=\begin{bmatrix}0.040,\,-0.509,\,-0.190,\,-0.030,\,0.162,\,-0.003,\,0.230,\,-0.218,\,-0.001,\,0.406\end{bmatrix}^{\top}\quad\text{(continuous outcome)}$.
Across both alignment regimes, the corrected two-stage procedure yields well-calibrated uncertainty for subgroup-specific treatment effects under the LR definition, with near-nominal empirical coverage across sample sizes and outcome types (Tables S3S4).
Table S3: Simulation results for the corrected estimator under a data generative mechanism where prognosis and treatment effect are nearly perfectly correlated.
@@
375 (125) UR 0.215 0.201 -0.120 0.499 0.955 0.082 0.081 -0.031 0.335 0.980
LR 0.213 0.210 -0.060 0.493 0.955 0.094 0.094 -0.017 0.335 0.970
750 (250) UR 0.110 0.107 -0.052 0.334 0.930 0.050 0.050 -0.005 0.229 0.950
LR 0.094 0.086 -0.088 0.334 0.950 0.042 0.042 -0.008 0.226 0.975
1500 (500) UR 0.057 0.051 -0.072 0.234 0.945 0.027 0.027 -0.019 0.155 0.940
LR 0.056 0.048 -0.091 0.232 0.940 0.022 0.022 -0.018 0.155 0.950
  • Abbreviations: UR, unlikely-responders; LR, likely-responders. $\text{n}_{\text{eval}}$: evaluation sample size; $\text{n}_{\text{design}}$: design-stage sample size.
  • Empirical correlation between $s(x)$ and $\Delta(x)$ in the full simulated population: binary outcome Spearman $\rho=0.66$, Pearson $r=0.84$; continuous outcome Spearman $\rho=1$, Pearson $r=1$.
Table S4: Simulation results for the corrected estimator under a data generative mechanism where prognosis and treatment effect are moderately negatively correlated.
MSE Var Bias $\overline{SE}$ Covrg MSE Var Bias $\overline{SE}$ Covrg
375 (125) UR 0.137 0.137 0.014 0.415 0.975 0.070 0.070 0.014 0.319 0.980
LR 0.170 0.169 -0.029 0.416 0.960 0.067 0.067 -0.012 0.321 0.990
750 (250) UR 0.068 0.068 0.018 0.288 0.985 0.035 0.035 0.005 0.217 0.970
LR 0.070 0.069 -0.028 0.287 0.985 0.030 0.030 -0.008 0.218 0.975
1500 (500) UR 0.034 0.033 0.012 0.201 0.975 0.017 0.017 -0.006 0.149 0.975
LR 0.030 0.030 -0.001 0.201 0.975 0.018 0.018 -0.014 0.149 0.975
  • Abbreviations: UR, unlikely-responders; LR, likely-responders. $\text{n}_{\text{eval}}$: evaluation sample size; $\text{n}_{\text{design}}$: design-stage sample size.
  • Empirical correlation between $s(x)$ and $\Delta(x)$ in the full simulated population: binary outcome Spearman $\rho=-0.17$, Pearson r=0.10r=-0.10; continuous outcome Spearman $\rho=-0.17$, Pearson r=0.18r=-0.18.
@@ D. Variable Importance in the BART Model for COMPILE COVID-19 Clinical Trial
In this section, we report results supplementary to Section 5.2 of the main manuscript, regarding the prognostic balancing score model developed in the design stage. In the context of BART, variable importance is conveyed through the Variable Inclusion Proportion (VIP).[2] The VIP represents the proportion of posterior MCMC samples in which a variable was included in a splitting rule in the sum-of-trees model, as described by Chipman et al.[7]. Figure S1 displays the VIPs from a BART model trained on all treated individuals. In order to sharpen our variable selection, we limited the model to 20 trees when computing the VIPs.
@@
Patients with preexisting conditions—cardiovascular disease, pulmonary disease, and diabetes—were, on average, predicted to have a lower probability of the outcome event compared to those without these conditions, despite all receiving CCP treatment. These findings align with results from the multivariate model reported by Wu et al.[30], which suggested that CCP treatment was more effective for patients with preexisting disease, given the reference levels of other pre-treatment characteristics. However, as shown in Figure 2 in main text, the contribution of these effects was minimal, making their clinical significance uncertain.